# LibMOON: A Gradient-based MultiObjective OptimizatioN Library in PyTorch

### v2

@@ Abstract
Keywords Mathematical tools  $\cdot$ Multiobjective optimization  $\cdot$ Pareto set learning  $\cdot$ Bayesian optimization
@@ 1 Introduction
MultiObjective Optimization problems (MOPs) are ubiquitous in machine learning. For instance, trustworthy machine learning (TML) includes fairness problems balancing the fairness level and the accuracy level [1, 2]; in robotics, it is necessary to balance several objectives (e.g., forward speed and energy consumption)  [3, 4]; similarly, recommendation systems face potentially conflicting objectives, such as novelty, accuracy, and diversity [5, 6, 7]. For all the applications above, the underlying optimization problem involves an MOP with $m$ objectives and can be (informally) defined as:
$\min_{\theta}{\bm{L}}({\bm{\theta}})=(L_{1}({\bm{\theta}}),\ldots,L_{m}({\bm{\theta}})),$ (1)
where $L_{1}({\bm{\theta}}),\ldots,L_{m}({\bm{\theta}})$ denote $m$ (potentially) conflicting objectives and we denote the size of the model parameter as $N\coloneqq|{\bm{\theta}}|$. Note that as informally defined above, Equation 1 is a vector optimization problem that does not necessarily admit a total ordering.
@@
For non-trivial multiobjective problems, no single solution can attain the minimum of all objectives simultaneously. To compare two solutions for an MOP, we introduce the concepts of dominance and Pareto optimality [8]. Dominance occurs when solution ${\bm{\theta}}^{(a)}$ satisfies $L_{i}({\bm{\theta}}^{(a)})\leq L_{i}({\bm{\theta}}^{(b)})$ for all $1\leq i\leq m$, with at least one strict inequality $L_{j}({\bm{\theta}}^{(a)})<L_{j}({\bm{\theta}}^{(b)})$. A solution is Pareto optimal if no other solution in the feasible region dominates it. The set of all Pareto optimal solutions is the Pareto set (PS), and its image is the Pareto front (PF).
Over the last few decades, multiobjective evolutionary algorithms (MOEAs) emerged as a widely used methodology for addressing MOPs. Due to their population-level designs, MOEAs can identify a diverse set of Pareto optimal solutions that approximate and represent the entire PF of a given MOP. Several libraries have emerged as standard platforms for fair comparison of MOEAs, including PlatEMO [9] (in Matlab supporting over 160 MOEAs), Pagmo [10] (in C++), and Pymoo [11] (in Python).
@@
Compared to MOEAs, gradient-based multiobjective optimization (MOO) methods are particularly suitable designed for large-scale machine learning tasks involving thousands to millions of neural network parameters. While these methods can find Pareto stationary solutions—solutions that cannot be locally improved in any objective—they are also effective approximators of Pareto solutions for nonconvex models, including deep neural networks [12].
In this paper, we roughly divide recent gradient-based MOO methods into two main groups. The first one focuses on finding a finite set of Pareto solutions to approximately represent the entire PF. Examples include Upper Bounded Multiple Gradient Descent Algorithm (MGDA-UB) [13] and Exact Pareto Optimization (EPO) [14]. The second, known as Pareto set learning (PSL) [15, 16, 17], develops a single neural model parameterized by ${\bm{\phi}}$ to represent the entire Pareto set. Since the parameter space of ${\bm{\phi}}$ is large, thus PSL is usually optimized by gradient-based methods.
@@
With the growing use of gradient-based MOO methods in neural networks [13, 18, 4, 16, 19, 20, 21], there is a pressing need for a standard library to benchmark comparisons and developments. To fulfill this demand, we introduce LibMOON, a gradient-based Multi-Objective Optimization Library supporting over twenty state-of-the-art methods.
We summarize our contributions as follows:
@@
    We propose the first modern gradient-based MOO library, called LibMOON. LibMOON is implemented under the PyTorch framework [22] and carefully designed to support GPU acceleration. LibMOON supports synthetic problems and a wide range of real-world multiobjective machine learning tasks (e.g., fair classification and multiobjective classification problems).
@@
    LibMOON supports over twenty state-of-the-art (SOTA) gradient-based MOO methods for constructing Pareto optimal solution sets, including those using finite solutions to approximate the whole PS/PF [14, 19], PSL methods [16, 15] aimed at approximating the entire Pareto set with a single neural network, and MOBO methods that designed to avoid frequent function evaluations.
@@
    LibMOON has already been open-sourced at https://github.com/xzhang2523/libmoon with document at https://libmoondocs.readthedocs.io/en/latest/, offering a reliable platform for comparing SOTA gradient-based MOO approaches and providing standard multiobjective testing benchmarks. Beyond examining its source code, LibMOON can be installed via pip install libmoon as an off-the-shelf gradient-based multiobjective tool.
@@
In this paper, bold letters represent vectors (e.g., ${\bm{\lambda}}$ for preference vectors, ${\bm{L}}()$ for vector objective functions), while non-bold letters represent scalars. The number of objectives is denoted by $m$, and $K$ represents the size of finite solution set. The preference vector ${\bm{\lambda}}$ lies in the $m$-dim simplex (${\bm{\Delta}}_{m}$), satisfying $\sum_{i=1}^{m}\lambda_{i}=1$ and $\lambda_{i}\geq 0$. The decision network parameter ${\bm{\theta}}$ has a size of $n$. Refer to Table 12 for the full notation table.
@@ 2 Related works
In this section, we review some methods for gradient-based multiobjective optimization in Section 2.1 and existing multiobjective optimization libraries in Section 2.2.
@@ 2.1 Gradient-based multiobjective optimization
Gradient-based MOO has a long research history. For instance, the famous convex optimization book [23][Chap 4] discusses how to use linear scalarization to convert MOO into a single-objective optimization (SOO) problem. However, in the last several decades, gradient-based methods were not a mainstream for MOO, with evolutionary algorithms (MOEAs) being more prevalent due to their ability to maintain a population, which is inherently suitable to approximate the Pareto set and avoid local optimal solutions. However, gradient-based MOO has gained growing popularity very recently with their applications in (deep) machine learning. A seminal work in this field is MGDA-UB [13], which introduced MOO concepts into deep learning by framing an MTL problem as an MOO. Since then, numerous approaches such as EPO [14], Pareto MultiTask learning (PMTL) [18], MOO with Stein Variational Gradient Descent (MOO-SVGD) [19], and methods for learning the entire Pareto set [15, 16, 24, 25, 26], have emerged. This paper categorized these methods by two main classes and implement these methods in a standard manner.
@@ 2.2 Multiobjective optimization libraries
A number of multiobjective libraries exist before our work. The difference between LibMOON and previous libraries is that LibMOON is designed to support gradient-based MOO methods. Comparisons with previous is provided in Table 1.
@@
LibMTL [27]. LibMTL is a Python-based library for multitask learning. The key difference between LibMOON and LibMTL is that LibMTL focuses on finding a single network to benefit all tasks, such as calculating a benign updating direction or optimizing a network architecture. In contrast, LibMOON addresses inherent trade-offs in machine learning problems, where improving one objective inevitably worsens others, and explores the distribution of Pareto solutions.
jMetal [28]. jMetal is a comprehensive Java framework designed for multi-objective optimization, leveraging the power of metaheuristics. It offers a highly flexible and extensible platform, making it accessible for users across various domains. With its user-friendly interface and robust architecture, jMetal has become a popular choice for researchers and practitioners in diverse application areas.
Pymoo [29]. Pymoo is a Python-based framework for multiobjective optimization problems with many local optima. Pymoo supports mainstream MOEA methods such as NSGA-II [30], NSGA-III [31, 32], MOEA/D [33], and SMS-EMOA [34]. Pymoo allows flexible algorithm customization with user-defined operators and includes tools for exploratory analysis and data visualization. It also offers multi-criteria decision-making strategies to help users select optimal solutions from the Pareto set.
PlatEMO [9]. PlatEMO is a MATLAB-based multiobjective optimization tool supporting over 160 evolutionary algorithms and a comprehensive set of test problems. It supports diverse problem types, including sparse, high-cost, large-scale, and multimodal. PlatEMO includes performance metrics for evaluating optimization efficacy and offers tools for visualization and interaction during the optimization process.
@@
Pagmo [10]. Pagmo, a C++ library, executes massively parallel multiobjective global optimization using efficient evolutionary algorithms and gradient-based techniques like simplex, SQP, and interior-point methods. It supports concurrent algorithm execution and inter-algorithm collaboration through asynchronous generalized island models, addressing a wide range of optimization challenges, including constrained, unconstrained, single-objective, multiobjective, continuous, integer, stochastic, and deterministic problems.
@@
EvoTorch [35] and EvoX [36]. EvoTorch accelerates evolutionary algorithms in PyTorch, while EvoX scales them to large, complex problems with GPU-accelerated parallel execution for single and multiobjective tasks, including synthetic and reinforcement learning. These libraries highlight a trend toward using GPU parallelization to manage larger problem scales and decision variables, which is also of the interest of this paper. This work focuses on creating a platform for efficient gradient-based multiobjective methods to solve large-scale machine learning problems in PyTorch.
Table 1: Comparison of different libraries and frameworks.
Name Language Publication Installable Key Feature
Pymoo Python 2020 Evolutionary computation (EC)
jMetal Java 2011 EC and meta-heuristics
PlatEMO Matlab 2017 × EC and meta-heuristics
Pagmo C++ 2020 × Global optimization
LibMTL Python 2023 A single solution benefiting all tasks
EvoTorch Python NA GPU acceleration for EC
EvoX Python 2024 GPU acceleration for EC
LibMOON Python NA GPU accelerated gradient-based MOO solvers
@@ 3 LibMOON: A gradient-based MOO library in PyTorch
This section introduces LibMOON . We first introduce its framework in Section 3.1, and briefly introducing its supporting problems and metrics. Then we introduce supported solvers in Sections 3.2, 3.3 and 3.4.
Figure 1: The Framework of LibMOON: LibMOON addresses synthetic, real-world and MTL problems with three categories of solvers: MOO, PSL and MOBO solvers.
@@ 3.1 Framework
Figure 1 demonstrates the framework of the proposed LibMOON library. LibMOON supports three categories of solvers: MOO solvers aiming to find a finite set of Pareto solutions satisfying certain requirements, PSL solvers aiming to learn whole PS with a single model, and MOBO solvers aiming to solve expensive problems. Each solver category is designed highly modulized and new solvers are easy to plugin LibMOON by rewriting only a small portion of code, e.g., the specific way of gradient manipulations  [^1] An example of adding a new method is given by https://libmoondocs.readthedocs.io/en/latest/develop/add_method.html.. MOO and PSL solvers support all synthetic, MTL, and realworld (RE) problems, while MOBO solvers support synthetic and RE problems.
Table 2: Supported multiobjective machine learning problems in LibMOON.
Method $L_{1}$ $L_{2}$
Fairness classification [37] Binary cross entropy DEO
Multiobjective classification [18] Cross entropy loss BR Cross entropy loss UL
MO machine learning [20] Mean square loss Mean square loss
MO distribution alignment Similarity 1 Similarity 2
  • DEO [38]: Difference of Equality of Opportunity.
@@
LibMOON supports a wide range of metrics namely, (1) hypervolume (HV), (2) inverted general distance (IGD), (3) fill distance (FD), (4) minimal distance ($\mathrm{l}_{\min}$), (5) smooth minimal distance ($\mathrm{sl}_{\min}$), (5) Spacing, (6) Span, (7) penalty-based intersection (PBI), (8) inner product (IP), (9) cross angle ($\vartheta$). Full descriptions of these indicators are provided in Table 13.
@@ 3.2 MOO solvers
Table 3: Supported finite Pareto solution solvers, their properties, and time complexity for one iteration.
Method Property Paper Complexity Pref. based
EPO [14] Exact solutions paper $O(m^{2}nK)$
HVGrad [39] Maximal HV solutions paper $O(m^{2}nK^{2})$
MGDA-UB [13] Arbitrary solutions paper $O(m^{2}nK)$
MOO-SVGD [19] Diverse solutions paper $O(m^{2}nK^{2})$
PMGDA [40] Specific solutions paper $O(m^{2}nK)$
PMTL [18] Sector-constrained solutions paper $O(m^{2}nK^{2})$
Random [41] Arbitrary solutions paper $O(m^{2}nK)$
Agg-LS [42] Convex hull solutions book $O(mnK)$
Agg-Tche [33] Exact solutions paper $O(mnK)$
Agg-mTche [43] Exact solutions paper $O(mnK)$
Agg-PBI [33] Approximate exact solutions paper $O(mnK)$
Agg-COSMOS [37] Approximate exact solutions paper $O(mnK)$
Agg-SmoothTche [25] Fast approximate exact solutions paper $O(mnK)$
  • $m$: number of objectives. $n$: number of decision variables. $K$: number of subproblems.
Aggregation function.

### v3

@@ Abstract
Keywords Mathematical tools  $\cdot$ Multiobjective optimization  $\cdot$ Pareto set learning  $\cdot$ Bayesian optimization  $\cdot$ Pareto Machine Learning
@@ 1 Introduction
MultiObjective Optimization problems (MOPs) are ubiquitous in machine learning. For instance, trustworthy machine learning includes algorithmic fairness problems balancing the fairness level and the accuracy level [1, 2]; in robotics, it is necessary to balance several objectives (e.g., forward speed and energy consumption)  [3, 4]; similarly, recommendation systems face potentially conflicting objectives, such as novelty, accuracy, and diversity [5, 6, 7]. For all the applications above, the underlying optimization problem involves an MOP with $m$ objectives and can be (informally) defined as:
$\min_{\bm{\theta}}{\bm{L}}({\bm{\theta}})=(L_{1}({\bm{\theta}}),\ldots,L_{m}({\bm{\theta}})),$ (1)
@@
where $L_{1}({\bm{\theta}}),\ldots,L_{m}({\bm{\theta}})$ denote $m$ (potentially) conflicting objectives and we denote the size of the model parameter as $N\coloneqq|{\bm{\theta}}|$. Note that as informally defined above, Equation 1 is a vector optimization problem that does not necessarily admit a total ordering. For a non-trivial MOP, no single solution can attain the minimum of all objectives simultaneously. To compare two solutions for an MOP, we introduce the concepts of dominance and Pareto optimality [8]. Dominance occurs when solution ${\bm{\theta}}^{(a)}$ satisfies $L_{i}({\bm{\theta}}^{(a)})\leq L_{i}({\bm{\theta}}^{(b)})$ for all $1\leq i\leq m$, with at least one strict inequality. A solution is Pareto optimal if no other solution in the feasible region dominates it. The set of all Pareto optimal solutions is called the Pareto set (PS), and its image set is called the Pareto front (PF).
@@
Over the last few decades, multiobjective evolutionary algorithms (MOEAs) emerged as a widely used methodology for addressing MOPs due to their population nature to approximate the PS. Several popular MOEA libraries have emerged, including PlatEMO (Matlab)  [9], Pagmo (C++) [10], and Pymoo (Python) [11]. Compared to MOEAs, gradient-based multiobjective optimization (MOO) methods are particularly suitable designed for large-scale machine learning tasks involving thousands to millions of neural network parameters. While, gradient based MOO methods can only find Pareto stationary solutions—solutions that cannot be locally improved, in practice Pareto stationary solutions approximate global Pareto solutions well.
@@
With the growing needs of gradient-based MOO methods for neural networks [12, 13, 4, 14], there is a pressing need for the development of a standard library to benchmark related algorithms and problems. For this reason, we introduce LibMOON, the first modern gradient-based MOO library supporting over twenty state-of-the-art (SOTA) methods. We summarize our contributions as:
@@
    We propose the first modern gradient-based MOO library, called LibMOON. LibMOON is implemented in PyTorch[15] and carefully designed to support GPU acceleration. LibMOON supports synthetic problems, real-world problems, and MO machine learning problems such as fair classification, MO classification, MO regression, MO distribution matching and etc.
@@
    LibMOON supports over twenty SOTA gradient-based MOO methods for constructing PS/PF, including those using finite solutions to approximate the whole PS/PF [16, 17], Pareto set learning (PSL) [18, 19] aimed at approximating the entire PS/PF with a single neural model, and MOBO methods that designed to avoid frequent function evaluations for
@@
    LibMOON has already been open-sourced at Github  [^1] https://github.com/xzhang2523/libmoon with document at LibMOON Docs   [^2] https://libmoondocs.readthedocs.io/en/latest/. Beyond examining its source code, LibMOON can be installed via pip install libmoon as an off-the-shelf gradient-based multiobjective tool for industrial usage.
@@
In this paper, bold letters represent vector (e.g., ${\bm{\lambda}}$ denotes a preference vector), while non-bold letters represent scalars. ${\bm{x}}_{k}$ denotes vector ${\bm{x}}$ at $k$-th iteration and $x_{k}$ denotes the $k$-th entry of ${\bm{x}}$. The preference vector ${\bm{\lambda}}$ lies in the $m$-dim simplex (${\bm{\Delta}}_{m}$), satisfying $\sum_{i=1}^{m}\lambda_{i}=1$ and $\lambda_{i}\geq 0$. The decision network parameter ${\bm{\theta}}$ has a size of $n$. For two $m$-D vectors ${\bm{x}}^{(a)}$ and ${\bm{x}}^{(b)}$, ${\bm{x}}^{(a)}\preceq{\bm{x}}^{(b)}$ means ${\bm{x}}_{i}^{(a)}\leq{\bm{x}}_{i}^{(b)}$ for all $i\in[m]$; ${\bm{x}}^{(b)}$, ${\bm{x}}^{(a)}\preceq_{\mathrm{strict}}{\bm{x}}^{(b)}$ means that ${\bm{x}}_{i}^{(a)}\leq{\bm{x}}_{i}^{(b)}$ for all $i\in[m]$ and for at least one index $j$, ${\bm{x}}_{j}^{(a)}<{\bm{x}}_{j}^{(b)}$. ${\bm{x}}^{(a)}\prec{\bm{x}}^{(b)}$ means that ${\bm{x}}_{i}^{(a)}<{\bm{x}}_{i}^{(b)}$ for all $i\in[m]$. Refer to Table 11 for the full notation meanings.
@@ 2.1 Gradient-based multiobjective optimization
Gradient-based MOO has a long-standing research history. For example, the well-known convex optimization [20][Chap 4] book outlines how linear scalarization can transform an MOO problem into a single-objective optimization (SOO) problem. However, for much of the past few decades, gradient-based methods have not been the primary approach for MOO, with MOEAs gaining more prominence due to their population-based approach, which is well-suited for approximating the PS and avoiding local optima. In recent years, however, gradient-based MOO has experienced a resurgence, particularly in (deep) machine learning, where these methods scale better with the number of decision variables. A pivotal contribution in this area is the MGDA-UB [12], which introduced MOO techniques into deep learning by casting multi-task learning (MTL) as an MOO problem. Since then, many approaches have followed, including EPO [16], Pareto Multi-Task Learning (PMTL) [13], MOO with Stein Variational Gradient Descent (MOO-SVGD) [17], and methods for learning the entire PS [19, 18, 21, 22, 23]. To make a fair comparison and for the ease of developing new methods, this paper implements them following a standardized manner.
@@ 2.2 Multiobjective optimization libraries
A number of multiobjective libraries exist before our work. LibMTL [24] is a library for multitask learning; the difference is that LibMOON is designed to study the distributions of Pareto solutions, while LibMTL is mainly designed to study how to find a single solution that benefits all objectives. The major difference between LibMOON and other previous libraries is LibMOON is a gradient-based library, while others are evolutionary computation-based libraries. Comparisons are provided in Table 1.
@@
LibMTL [24] is a Python-based library for multitask learning. LibMTL aims to find a single network to benefit all tasks, such as calculating a benign updating direction or optimizing a network architecture. In contrast, LibMOON addresses inherent trade-offs in machine learning problems, where improving one objective inevitably worsens others, and explores the distribution of Pareto solutions.
jMetal [25], Pymoo [26] and PlatEMO [9] are Java, python and Matlab frameworks for MOEAs, supporting popular methods such as NSGA-III [27, 28], MOEA/D [29], and SMS-EMOA [30]. Pymoo allows flexible algorithm customization with user-defined operators and data visualization. PlatEMO is a MATLAB-based multiobjective optimization tool supporting over 160 MOEAs and a comprehensive test problems, including sparse, high-cost, large-scale, and multimodal. PlatEMO also contains a number of metrics and supporting visualization during the optimization process.
@@
Pagmo [10] is a C++ library for parallel multiobjective global optimization, utilizing evolutionary algorithms and gradient-based methods like simplex, SQP, and interior-point techniques. It supports constrained, unconstrained, single- and multi-objective, continuous, integer, stochastic, and deterministic optimization problems.
Table 1: Previous MOO libraries and LibMOON.
Name Language Year Key Features
Pymoo Python 2020
(1) Evolutionary computation (EC)
(2) Zero-order methods
(3) Diverse problem types
jMetal Java 2011
(1) Single-/multi-objective optimization
(2) Parallel algorithms
(3) Diverse problem types
PlatEMO Matlab 2017
(1) Over 160 MOEAs
(2) Various figure demonstrations
(3) Powerful and friendly GUI
Pagmo C++ 2020
(1) Global optimization
(2) Parallel optimization
LibMTL Python 2023
(1) Unified codebase
(2) Comprehensive SOTA MTL methods
(3) Flexible extension for new methods
EvoTorch Python N/A
(1) Distribution-based search algorithms
(2) Population-based search algorithms
(3) Multiple CPUs, GPUs, computers
EvoX Python 2024
(1) GPU acceleration optimization
(2) Single-/multi-objective optimization
(3) Neuroevolution/RL tasks
LibMOON Python 2024
(1) GPU-accelerated gradient solvers
(2) Pareto set learners
(3) Large-scale (millions # params.) ML tasks
@@
EvoTorch [31] and EvoX [32]. EvoTorch accelerates evolutionary algorithms in PyTorch, while EvoX scales them to large, complex problems with GPU-accelerated parallel execution for single and multiobjective tasks, including synthetic problems and reinforcement learning.
@@ 3 LibMOON: A gradient-based MOO library in PyTorch
This section introduces LibMOON. We introduce its framework in Section 3.1, and briefly introducing its supporting problems and metrics. Then we introduce supported solvers in Sections 3.2, 3.3 and 3.4.
Figure 1: Supported solvers and problems in LibMOON: LibMOON addresses synthetic, real-world and MTL problems with three categories of solvers: MOO, PSL and MOBO solvers.
@@ 3.1 Framework
Figure 1 demonstrates LibMOON components, including three categories of solvers: MOO solvers aiming to find a finite set of Pareto solutions satisfying certain requirements, Pareto set learning (PSL) solvers aiming to learn whole PS with a single model, and MOBO solvers aiming to solve expensive MO problems. Each solver category is designed highly modulized and new solvers are easy to plugin LibMOON by rewriting only a small portion of code, e.g., the specific algorithm of gradient manipulations  [^3] An example of adding a new solver is provided in the LibMOON Doc.. MOO and PSL solvers support all synthetic, MTL, and realworld (RE) problems, while MOBO solvers support synthetic and RE problems.
@@
Table 2: Supported MO machine learning problems.
Method $L_{1}$ $L_{2}$
Fairness classification [33] BCE DEO
MO classification [13] CE - BR CE - UL
MO regression [34] MSE MSE
MO distribution matching $D(\cdot\|\cdot)$ $D(\cdot\|\cdot)$
  • BCE: Binary Cross Entropy; DEO [35]: Difference of Equality of Opportunity; CE: Cross Entropy; BR: Bottom Right; UL: Upper Left; MSE: Mean Square Error.
@@
LibMOON supports a number of metrics, (1) hypervolume (HV), (2) inverted general distance (IGD), (3) fill distance (FD), (4) minimal distance ($\mathrm{l}_{\min}$), (5) smooth minimal distance ($\mathrm{sl}_{\min}$), (5) Spacing, (6) Span, (7) penalty-based intersection (PBI), (8) inner product (IP), (9) cross angle ($\vartheta$). Full descriptions of these indicators are provided in Section A.3.
@@ 3.2 MOO solvers
In this paper, we specially mean MOO solvers are solvers to find a set of Pareto solutions. The most simple and commonly used method is convert the MOO problem to a single objective optimization problem through some aggregation functions.
Aggregation-based methods.